OpenStudy (anonymous):

What is the exact value of cos(7π/12)+cos(π/ 12)? 0 2√ 4 2√ 2 −2√ /2

OpenStudy (anonymous):

use the formula\[\cos a+\cos b=\frac{1}{2} \cos \left(\frac{a+b}{2}\right) \cos \left(\frac{a-b}{2}\right)\]

OpenStudy (anonymous):

i got the wrong answer i think i did it wrong i still dont understand

OpenStudy (anonymous):

show me your steps please

OpenStudy (anonymous):

i figured it out, i just did my steps wrong thank you

OpenStudy (anonymous):

\[ 2 \cos \left(\frac{1}{2} \left(\frac{7 \pi }{12}+\frac{\pi }{12}\right)\right) \cos \left(\frac{1}{2} \left(\frac{7 \pi }{12}-\frac{\pi }{12}\right)\right)=\frac{1}{\sqrt{2}} \] I did not get any choices. Have I done something wrong?

OpenStudy (anonymous):

i got \[\frac{\sqrt{2}}{2}\]too !!

OpenStudy (anonymous):

thats the answer!!